OpenStudy (anonymous):

I am clueless and the book is no help. (sine of arccosine equation) sin(2(cos^-1)√2/2) please walk me through it, just the answer doesnt help.

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

does it help to know that \[\cos^{-1}(\frac{\sqrt2}{2})=\frac{\pi}{4}\]

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

it should

OpenStudy (misty1212):

then \[2\times \frac{\pi}{4}=\frac{\pi}{2}\]

OpenStudy (misty1212):

final answer is what you get when you find \[\sin(\frac{\pi}{2})\]

OpenStudy (anonymous):

so it would be the sine of pi/2?

OpenStudy (misty1212):

yup

OpenStudy (anonymous):

okay, Thanks very VERY much!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]